Jamey planted a sapling in his back yard for his school's ecology program. The sapling was ten inches high when he planted it. Two years later it was one foot and ten inches high. What is the percent of change in the tree's growth during this time? Label your answer as an increase or decrease.
Answer by stanbon(75887)

Jamey planted a sapling in his back yard for his school's ecology program. The sapling was ten inches high when he planted it. Two years later it was one foot and ten inches high. What is the percent of change in the tree's growth during this time? Label your answer as an increase or decrease.
------
Quantity increase:
start: 10 inches
after 2 years: 12+10 = 22 inches
Quantity increase = 22-10 = 12 inches
---
Percent increase: 12/10 = 1.2 = 120%
========================================
cheers,
Stan H.